(y−1)dx−(x−y−1)dy=0



1
Expert's answer
2021-09-28T01:34:59-0400

"(y-1)dx-(x-y-1)dy=0\\
M=y-1 \implies M_y=1\\
N=-(x-y-1) \implies N_x=-1\\
\text{Thus } M_y\neq N_x \implies \text{ It is not exact}\\
\text{To get the integrating factor;}\\
\frac{N_x-M_y}{M}=-\frac{2}{y-1}. \\
I.F=e^{\int -\frac{2}{y-1}dy}=e^{-2\ln (y-1)}=(y-1)^{-2}\\
\text{Multiply the DE by I.F}\\
(y-1)^{-1}dx-\frac{x-y-1}{(y-1)^2}dy=0\\
M_y=N_x=-\frac{1}{(y-1)^2}\\
F_x=(y-1)^{-1}dx\\
F=\int(y-1)^{-1}dx\\
F=x(y-1)^{-1}+\Phi(y)\\
\text{Differentiate w.r.t y}\\
F_y=-\frac{x}{(y-1)^2}+\Phi'(y)\\
-\frac{x-y-1}{(y-1)^2}=-\frac{x}{(y-1)^2}+\Phi'(y)\\
\Phi'(y)=-\frac{y+1}{(y-1)^2}\\
\Phi(y)=\frac{2}{y-1}-\ln(y-1)\\
\text{Hence}\\
F=x(y-1)^{-1}+\frac{2}{y-1}-\ln(y-1)"